XYZ Ltd last paid a dividend of $0.20 three years ago. Today the company announced they will resume paying dividends. The planned dividends are $0.65 in one year's time, $0.75 in two years' time, and thereafter dividends will increase by a constant rate of 3% p.a. indefinitely. If the required rate of return for XYZ is 12%, what is a fair price for one share today?
Answer #1

Step 1: Computation of market price at the end of year 2 using Gordon Growth Model

Using Gordon Growth Model

P2 = D3 / (Ke – g)

Where,

P2 - Market price at the end of year 2 = ?

D3 - Expected dividend in year 3 = D2*(1+g) = .75*1.03 = 0.7725

Ke – Cost of equity = 12%

G – Growth rate in dividend = 3%

P2 = 0.7725/(.12-.03)

= 0.7725/.09

= $8.5833

Step 2: Computing current share price by discounting the cashflow at required return

Year Dividend PVF@12% Present Value (Cashflow*PVF)
1 0.6500                   0.893 0.58
2 9.3333 (.75+8.5833)                   0.797 7.44

Current share price = \sumCashflow*PVF

= .58+7.44

= $8.02

You can use the equation 1/(1+i)^n to find PVF using calculator